### v1

# Is the Cure Still Worse Than the Disease? Test Overfitting by LLMs in Automated Program Repair

Abstract.
Automated program repair has been shown to be susceptible to generating repaired code that passes on seen tests but fails on a hold-out set of hidden tests. This problem, dubbed test overfitting, has been identified and studied before the rise of large language models. We experimentally study how much test overfitting is still a problem today, using repository-level SWE-bench tasks.
LLMs, SWE Patches, Reproduction Tests
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Automated program repair (APR), which tries to automatically fix bugs in code, has long been an active research topic in the software engineering community (legoues_pradel_roychoudhury_2019). Thanks to SWE-bench (jimenez_et_al_2024), a recent benchmark for issue-driven repository-level APR, it is now also a hot research topic in the artificial intelligence (AI) community. Most solutions use large language models (LLMs) in prompted workflows without fine-tuning, such as SWE-Agent (yang_et_al_2024), Agentless (xia_et_al_2025), or CodeMonkeys (ehrlich_et_al_2025). SWE-bench evaluates a candidate code repair using hidden (black-box) tests. Therefore, solutions often leverage a disjoint set of seen (white-box) tests to select or improve a good candidate code repair (yang_et_al_2024; xia_et_al_2025; ehrlich_et_al_2025). Unfortunately, doing so creates the risk of test overfitting, where the code generated by APR passes white-box tests but fails black-box tests. Even though Smith et al. identified the test overfitting problem ten years ago (smith_et_al_2015), current LLM-driven APR solutions do not account for it.
Therefore, this paper sets out to answer the question: is test overfitting still a problem in the era of LLM-based APR workflows? Besides Smith et al.’s work (smith_et_al_2015) which predates LLMs, the closest related work on test overfitting is by Stroebl et al. (stroebl_kapoor_narayanan_2024). However, unlike our work, it does not use repository-level benchmarks, and uses white-box tests only to select one among a set of candidate code repairs, not to improve candidate code repairs. Some recent workflows have sophisticated loops of improving code repairs along with white-box tests (ehrlich_et_al_2025; li_et_al_2025_sstar). This approach might exacerbate test overfitting, which the papers unfortunately do not explore.
In contrast to the prior work, our paper directly investigates test overfitting in repository-level tasks with state-of-the-art (SoTA) LLM-based workflows. It starts from an initial candidate code generated by the SoTA APR solution Agentless (xia_et_al_2025) and an initial candidate test generated by the SoTA issue reproduction solution e-Otter++ (ahmed_et_al_2025_eotter). This paper compares the behavior of APR code on both white-box and black-box tests to answer the following research questions: RQ1. Can LLMs overfit on white-box tests? (Yes, they can and do.) RQ2. How does improving code based on white-box tests affect test overfitting? (Overfitting goes up for affected instances.) RQ3. How much would revealing the black-box tests increase apparent resolution rate as measured by the black-box tests? (Not vastly more than using white-box tests.) The experiments are based on the TDD-bench Verified (ahmed_et_al_2024) benchmark for test-driven development and use two LLMs, Claude-3.7 Sonnet and GPT-4o. This paper makes the following contributions.
  • It measures the degree of test overfitting by LLMs in repository-level APR tasks.
  • It describes techniques to reduce overfitting and reports how much they help overfitting and hurt performance.
  • It also includes a limit study of how much an oracle with access to the golden black-box tests could help performance.
Our paper title is inspired by Smith et al. (smith_et_al_2015). If the disease is code bugs and the cure is APR with white-box tests, Smith et al. found that the cure was indeed worse than the disease: APR decreased performance on black-box tests. Our findings are less drastic: on average across all benchmark instances, APR with white-box tests improves performance on black-box tests. But there is also still test overfitting to beware of—the cure can still cause harm.
2. Background and Related Work
Test overfitting happens when an APR workflow generates code that passes white-box tests but fails held-out black-box tests (smith_et_al_2015). Test overfitting is similar to overfitting a model to its training data, where the model performs worse on held-out test data; but note that in this work, we use only pre-trained LLMs without fine-tuning. Instead, test overfitting means that generated repaired code is too narrow to generalize to held-out black-box tests. Stroebl et al. report that this happens when a loop repeatedly generates code for the HumanEval or MBPP benchmarks until finding code that passes white-box tests (stroebl_kapoor_narayanan_2024). Unlike Stroebl et al., we experiment with repository-level benchmarks and we also explore workflows that improve code based on tests.
SWE-bench is an APR benchmark where the input comprises an issue and the original code of an entire repository before issue resolution and the output is repaired code that must pass a held-out test set (jimenez_et_al_2024). Those held-out tests are unavailable, hidden from the APR workflow. SWE-bench Verified is a subset of 500 high-quality SWE-bench instances, targeted by most LLM-based APR workflows (chowdhury_et_al_2024). SWT-bench (mundler_et_al_2024) and TDD-Bench Verified (ahmed_et_al_2024) are reproduction test benchmarks derived from SWE-bench and SWE-Bench Verified. A generated output test must fail on the original code to reproduce the issue, and must pass on the hidden ground-truth code patch to confirm that it resolves the issue (mundler_et_al_2024).
Several LLM-based workflows leverage white-box tests. CodeT generates many code candidates and many candidate white-box tests, then uses a dual execution agreement to pick a single output code (chen_et_al_2023). Agentless generates up to 40 candidate codes for SWE-bench and generates a white-box reproduction test, then uses the test to filter the candidate codes as part of code selection (xia_et_al_2025). Both S* (li_et_al_2025_sstar) and CodeMonkeys (ehrlich_et_al_2025) also generate both candidate codes and candidate tests, followed by a loop where the LLM improves the code based on the test and vice versa. These papers might suffer from test overfitting but do not study it.
Test overfitting is related to reward hacking, where a reward is a reinforcement-learning (RL) objective, and a hack finds a loop-hole for improving the reward while circumventing its intention. CURE uses RL to co-evolve an LLM to generate both code and white-box tests, and might be susceptible to test overfitting and reward hacking (wang_et_al_2025). Baker et al. found that RL to improve frontier LLMs for repository-level APR tasks can cause the reward hack of disabling tests instead of repairing code (baker_et_al_2025).
3. Methodology
Let $c_{\mathrm{old}}$ be the original code in the repository before a given issue is fixed. This work uses Agentless (xia_et_al_2025) to generate new code $c_{\mathrm{new}}$ for APR and e-Otter++ (ahmed_et_al_2025_eotter) to generate a white-box test $t$ for issue reproduction. Being generated by LLMs, both $c_{\mathrm{new}}$ and $t$ are possibly imperfect. We run the generated test $t$ on the generated code $c_{\mathrm{new}}$. If it fails, we use a test-based code improvement approach to generate a better $c_{\mathrm{new}}$ that passes the (possibly also improved) white-box test $t$. This section briefly discusses the generation of the initial $c_{\mathrm{new}}$ and $t$ and our code improvement approach.
3.1. Code Patch Generation: Agentless
Agentless performs hierarchical localization followed by repair and patch validation (xia_et_al_2025). Localization starts by identifying suspicious files using both LLM prompting and embedding-based retrieval, filtering out irrelevant files. It then narrows down to relevant classes or functions using a concise skeleton representation of files. Finally, it pinpoints exact edit locations at the line level. Repair generates multiple candidates for new code $c_{\textrm{new}}$ in a lightweight search-replace diff format rather than rewriting entire code blocks, which minimizes hallucinations and enhances cost-efficiency. Patch validation automatically generates reproduction tests $t$ (when none are provided) and combines them with existing regression tests in the repository to select a single $c_{\textrm{new}}$ via majority voting. This modular structure allows Agentless to remain interpretable, easily debuggable, and effective without relying on complex environment interaction or planning mechanisms typical of agents. With the Claude-3.7-Sonnet model, Agentless achieves a 50.8% issue resolution rate on SWE-bench Verified.
3.2. Reproduction Test Generation: e-Otter++
e-Otter++ (ahmed_et_al_2025_eotter)is a system designed to generate reproduction tests $t$ from software issue descriptions, such as bug reports or feature requests, before the corresponding code patch exists. It supports test-driven development workflows and helps validate code patches produced by APR workflows. The system augments LLM outputs with checks and repairs, such as linter checks and fixing import statements. It also employs a self-reflective action planner that iteratively decides which parts of the code to ‘read’ and where and how to write or modify test code. After generating candidate tests, e-Otter++ obtains execution feedback, such as runtime errors or successes, and uses this information to refine prompts and guide test selection. Additionally, e-Otter++ leverages issue morphing–transforming a bug report into multiple variants to generate diverse tests. It then uses a ranker that combines running on imperfect surrogate code patches and coverage to select a single top-ranked test $t$. As a result, e-Otter++ achieves a fail-to-pass success rate of 63% on the TDD-Bench Verified benchmark using Claude-3.7-Sonnet.
3.3. Test-based Code Refinement
Figure 1. Overview of test-based code improvement.
Figure 1 shows our test-based code refinement approach. We start with the initial candidate code $c_{\textrm{new}}$ from Agentless and the initial candidate test $t$ from e-Otter++. We execute the test on $c_{\textrm{old}}$ and $c_{\textrm{new}}$. If the test goes from fail-to-pass, we immediately stop and declare success. Otherwise, we collect the execution logs from both test executions, the focal functions modified by $c_{\textrm{new}}$, and the test function $t$ and provide them to an LLM-based critic. We ask the critic which part to modify (focal or test). Inspired by e-Otter++, we also collect additional information such as the buggy line, relevant issue line, and lookup function (the function the model wants to see to fix it). Then, we ask the LLM to modify the focal or test, exposing the counterpart (e.g., for focal, the white-box test) and additional information. After updating the focal or test, if the test is still not fail-to-pass, we need to decide whether to stick with our initial focal or test or update it with the recently generated focal or test, even though the test is not fail-to-pass. To do this, we use a reward function to guide our decision: $\mathit{Reward}(c_{\mathit{old}},c_{\mathit{new}},t)\,=$
$\frac{1}{3}\mathit{isFail}(t,c_{\mathrm{old}})+\frac{1}{3}\mathit{isPass}(t,c_{\mathrm{new}})+\frac{1}{3}\mathit{coverage}(c_{\mathit{old}},c_{\mathit{new}},t)$
In the reward function, the first two components are binary, but coverage can vary from 0 to 1. Since we know both the patch and the test, we can compute how many deleted lines in $c_{\mathrm{old}}$ and how many added lines in $c_{\mathrm{new}}$ have been covered by the test. We divide the covered lines by the total number of added or deleted lines to compute coverage. If our new code+test pair achieves a higher score than the previous one, we replace it with the new one. We repeat the improvement loop for at most 15 iterations if the white-box test does not change from fail to pass.
4. Results
4.1. Expermental Setup
The experiments in this paper are based on Agentless [^1] https://github.com/OpenAutoCoder/Agentless-generated code patches and e-Otter++ [^2] https://zenodo.org/records/16755224-generated tests. Agentless was evaluated on SWE-bench Verified, and e-Otter++ was evaluated on TDD-Bench Verified, a subset of 449 instances from SWE-bench Verified. We evaluated on all 449 instances. The experiments use two frontier models: Claude-3.7-Sonnet and GPT-4o. Our primary metric is test overfitting rate. A code patch is overfitting tests if it passes white-box tests but fails black-box tests. Dividing the number of instances where that happes by the total number of instances yields the overfitting rate.
Figure 2. How severe is test overfitting? (Claude-3.7-Sonnet)
Table 1. Test overfitting rate of automatic program repair.
Model Setup
# of
Sample
Resolved Unresolved
%of
Overfitting
Claude-3.7-Sonnet w/o Refinement 229 179 50 21.8
w/ Refinement 251 187 64 25.5
GPT-4o w/o Refinement 176 118 58 33.0
w/ Refinement 198 127 71 35.9
4.2. LLMs Overfit on White-box Tests (RQ1)
To answer this research question, we run the e-Otter++ generated white-box tests on Agentless patches and see how many of them go from fail-to-pass. Figure 2 shows that with the Claude-3.7-Sonnet model, 229 tests actually pass on the initial patches. Does this mean that the black-box golden tests will pass on these patches? The answer is no. There are 50 instances where the black-box test could not pass, signaling potential overfitting (50/229=21.8%). For the GPT-4o model, the overfitting rate is 33.0%, even higher than the Claude-3.7-Sonnet model (see Table 1).
4.3. Overfitting with Code Refinement (RQ2)
Tests have been used to validate SWE-patches. What about using tests to improve patches at generation time? Using our test-based code refinement technique, we attempt to generate an improved code+test pair for each of the 220 code+test pairs where the tests did not initially go from fail to pass. Refinement succeeded in generating 22 pairs where the white-box test goes from fail to pass. Unfortunately, 14 out of the 22 fail black-box testing, increasing the overfitting. If we include these in our existing overfitted samples, the overfitting rate increases from 21.8% to 25.5% (Table 1). GPT-4o experiences a similar increase in overfitting, indicating that it is difficult to improve code patches by exposing tests to LLMs.
In the above experiments, the refinement loop exposed the white-box tests and additional information to the LLMs. One idea to reduce overfitting might be to hide the test and additional information, so the LLMs may not be able to trick the test. Indeed, hiding the test causes a drop in agreement between code patches and white-box tests. However, the number of instances passing black-box tests also goes down a bit. If we remove the additional information, it becomes more difficult for LLMs to overfit the patch. Unfortunately, we still see a high overfitting rate even after this intuitive overfitting reduction approach. The model can find a way by knowing that the existing test is not fail-to-pass (third row in Table 2). Note that this will not have any impact on instances that were already overfitted before the refinement.
Table 2. Test overfitting rate of repaired samples.
Model
Expose
Test
Augment
Prompt
# of
Sample
Resolved Unresolved
%of
Overfitting
Claude-3.7 Sonnet 22 7 15 68.2
18 5 13 72.2
14 6 8 57.1
GPT-4o 22 9 13 59.1
17 5 12 70.6
16 5 11 68.8
Table 2 shows that exposing tests and prompt augmentation resolved 7 samples. Do all these 7 samples help us resolve additional instances? The answer is no. The initial code patch for one instance was already passing on the black-box test. And besides these 7, there were 2 instances that went from resolved to non-resolved. So the net gain is +4 instead of +7. That indicates the actual gain from test-based refinement is less than the apparent gain.
4.4. Impact of Revealing Black-box Tests (RQ3)
Table 3 shows how much overfitting occurs when using some of the golden tests for code selection and repair. For this experiment, we divide black-box tests into two groups: fail-to-pass reproduction tests and pass-to-pass regression tests. We only reveal the reproduction tests (so they are no longer black-box) but continue to hide the regression tests. This is a limit study, because in practice, all black-box tests are hidden. The overfitting rate, which means passing now-revealed reproduction tests but failing still-hidden regression tests, is low (13/223=5.8% and 22/194=11.3% with Claude and GPT-4o, respectively). Apart from overfitting, this also shows how well the model performs when some black-box tests are exposed. When the model passes the golden reproduction tests, the probability that the issue is resolved is high, but not 100%, because regression tests can still fail. The improvement from using black-box tests in issue resolution is modest: the number of resolved instances with the Claude model goes up from 229 to 243, which is around a 3% increase. Overfitting in this setup is much less since we are evaluating against the golden tests–however, the golden tests themselves are really just a proxy for “correctness”, against which we might still be overfitting (but we do not have a good way to measure this).
Table 3. Impact of revealing golden test on issue resolution.
Model
# of
Sample
Resolved Unresolved
%of
Overfitting
Claude-3.7 Sonnet 223 210 13 5.8
GPT-4o 194 172 22 11.3
5. Discussion
5.1. LLMs’ Preference on Focal and Test
Test-based code refinement gives LLMs a choice to modify both the code (focal function) and/or the test. Table 4 shows that for instances where refinement succeeds, the model mostly elects to modify the focal function instead of the test. For each scenario, the number of test updates is very low compared to focal updates. Also, there are few cases where we see the LLMs try to fix both the focal function and the test. That indicates the model has a preference for focal modification and believes that the tests are mostly perfect.
Table 4. LLMs preference for focal/test modifications
Model
Expose
Test
Prompt
Augment
# of Sample
Focal
Update
Test
Update
Both
Update
Claude-3.7 Sonnet 22 18 3 1
18 15 3 0
14 13 0 1
(Gold) (Gold) 25 23 0 2
GPT-4o 22 16 2 4
17 13 1 3
16 11 1 4
(Gold) (Gold) 28 20 2 6
Figure 3. Coverage of the unbiased and overfitted patches.
5.2. Coverage for Overfitted Sample
Figure 3 compares the coverage between unbiased code patches (both white-box and black-box tests pass) and overfitted patches for Claude-3.7. Note that the tests are generated by e-Otter++ and the code patches are from Agentless. The median for unbiased patches is 1, while for overfitted patches it is less than 0.8. Also, the mean coverage is higher for unbiased patches (0.95 vs. 0.9). These results indicate that even though both unbiased and overfitted patches pass the white-box test, overfitted patches have lower coverage.
5.3. What Could Happen with Overfitting?
Frontier LLMs are capable of tricking the test and changing the code to narrowly passes that particular test. For example, instance id “astropy__astropy-14598” complains about an inconsistency in double and single quotes. The issue writer notes that while dealing with a null string, the double quotes become single quotes. To trick the white-box test, LLMs modify their initial solution (see Figure 4(a)) to a solution where “replace” is invoked three times just to pass the white-box test (see Figure 4(b)), which is clearly an indication of LLMs overfiting. We have even seen LLMs using reflection to access private methods to pass the white-box test, creating a security risk.
(a) Initial code patch
(b) Overfitted refined code patch
Figure 4. Example where test-based refinement overfits.
6. Threats to Validity
One of the main limitations of our work is that the benchmarks include only Python repositories. Therefore, our findings may not generalize to other programming languages. We experimented with two models but did not investigate model contamination or memorization. However, prior works show that model-generated tests are very different from those written by developers. On another note, our approach may be considered a way to measure contamination by exposing overfitted patches. We leave this for future research. Other minor limitations include our reliance on the coverage package for reward formulation. The Python coverage package sometimes does not work due to dependency issues or multithreading. However, this happens in less than 1% of cases, and we assign a coverage of 1 to mitigate the impact of imperfect coverage reports. Our results can also be affected by test flakiness. However, we use pre-built docker containers for the experiments, which should minimize the impact of flakiness.
7. Conclusion
In this work, we systematically show that test overfitting is still a problem in LLM-based program repair. We also show that test-based refinement may resolve some additional issues but increases overfitting. Therefore, using tests for patch generation needs more deliberation. We share the initial and modified code patches at this link: https://zenodo.org/records/17227319
8. Future Plans
This paper demonstrates that overfitting exists and shows attempts to minimize it in a test-based code repair system. However, we have not proposed any methods to address overfitting in cases where the initial white-box test changes from fail to pass on overfitted patches. This is an important problem because the majority of overfitted patches come from this category. In the future, we will try to design an approach to automatically detect these overfitted patches and propose a solution to address the issue.

### v2

@@
-Is the Cure Still Worse Than the Disease? Test Overfitting by LLMs in Automated Program Repair
+Investigating Test Overfitting on SWE-bench
 Abstract.
-Automated program repair has been shown to be susceptible to generating repaired code that passes on seen tests but fails on a hold-out set of hidden tests.
-This problem, dubbed test overfitting, has been identified and studied before the rise of large language models.
-We experimentally study how much test overfitting is still a problem today, using repository-level SWE-bench tasks.
+Tests can be useful towards resolving issues on code repositories.
+However, relying too much on tests for issue resolution can lead to code that technically passes observed tests but actually misses important cases or even breaks functionality.
+This problem, called test overfitting, is exacerbated by the fact that issues usually lack readily executable tests.
+Instead, several issue resolution systems use tests auto-generated from issues, which may be imperfect.
+Some systems even iteratively refine code and tests jointly.
+This paper presents the first empirical study of test overfitting in this setting.
 LLMs, SWE Patches, Reproduction Tests
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
 1. Introduction
-Automated program repair (APR), which tries to automatically fix bugs in code, has long been an active research topic in the software engineering community (legoues_pradel_roychoudhury_2019).
-Thanks to SWE-bench (jimenez_et_al_2024), a recent benchmark for issue-driven repository-level APR, it is now also a hot research topic in the artificial intelligence (AI) community.
-Most solutions use large language models (LLMs) in prompted workflows without fine-tuning, such as SWE-Agent (yang_et_al_2024), Agentless (xia_et_al_2025), or CodeMonkeys (ehrlich_et_al_2025).
-SWE-bench evaluates a candidate code repair using hidden (black-box) tests.
-Therefore, solutions often leverage a disjoint set of seen (white-box) tests to select or improve a good candidate code repair (yang_et_al_2024; xia_et_al_2025; ehrlich_et_al_2025).
-Unfortunately, doing so creates the risk of test overfitting, where the code generated by APR passes white-box tests but fails black-box tests.
-Even though Smith et al.
-identified the test overfitting problem ten years ago (smith_et_al_2015), current LLM-driven APR solutions do not account for it.
-Therefore, this paper sets out to answer the question: is test overfitting still a problem in the era of LLM-based APR workflows?
-Besides Smith et al.’s work (smith_et_al_2015) which predates LLMs, the closest related work on test overfitting is by Stroebl et al.
-(stroebl_kapoor_narayanan_2024).
-However, unlike our work, it does not use repository-level benchmarks, and uses white-box tests only to select one among a set of candidate code repairs, not to improve candidate code repairs.
-Some recent workflows have sophisticated loops of improving code repairs along with white-box tests (ehrlich_et_al_2025; li_et_al_2025_sstar).
-This approach might exacerbate test overfitting, which the papers unfortunately do not explore.
-In contrast to the prior work, our paper directly investigates test overfitting in repository-level tasks with state-of-the-art (SoTA) LLM-based workflows.
-It starts from an initial candidate code generated by the SoTA APR solution Agentless (xia_et_al_2025) and an initial candidate test generated by the SoTA issue reproduction solution e-Otter++ (ahmed_et_al_2025_eotter).
-This paper compares the behavior of APR code on both white-box and black-box tests to answer the following research questions: RQ1.
-Can LLMs overfit on white-box tests?
-(Yes, they can and do.) RQ2.
-How does improving code based on white-box tests affect test overfitting?
-(Overfitting goes up for affected instances.) RQ3.
-How much would revealing the black-box tests increase apparent resolution rate as measured by the black-box tests?
-(Not vastly more than using white-box tests.) The experiments are based on the TDD-bench Verified (ahmed_et_al_2024) benchmark for test-driven development and use two LLMs, Claude-3.7 Sonnet and GPT-4o.
+Issue resolution systems, which automatically resolve bug reports or feature requests in code repositories, are useful.
+They are also challenging to build, because issue descriptions tend to be informal and code repositories are often large and complex (Jimenez et al., 2024).
+Several issue resolution systems leverage tests: for code candidate selection in parallel inference scaling (Arora et al., 2024; Gao et al., 2025; Jain et al., 2025; Ruan et al., 2024; Xia et al., 2025; Zhang et al., 2025), code refinement in sequential inference scaling (Chen et al., 2024; Li et al., 2025b; Yang et al., 2024; Zhang et al., 2024), or both (Ehrlich et al., 2025; Li et al., 2025c).
+These systems rely upon the implicit premise that tests offer a reliable signal in this setting.
+Unfortunately, little is known about whether this premise holds.
+To the contrary, startling results in adjacent settings cast doubt upon it (Baker et al., 2025; Smith et al., 2015; Stroebl et al., 2024).
+Relying upon tests for issue resolution can cause test overfitting: generated code that narrowly passes the observed tests but breaks other functionality.
+This paper empirically investigates the problem of test overfitting in repository-level issue resolution, on benchmarks such as SWE-bench (Jimenez et al., 2024).
+An issue resolution system takes an issue description $d_{\textrm{issue}}$ along with the original code $c_{\textrm{old}}$ of a repository as input and produces modified code $c_{\textrm{new}}$ as output.
+While many repositories come with regression tests $t_{\textrm{old}}\subset c_{\textrm{old}}$, those only test existing behavior, not the desired functionality of $d_{\textrm{issue}}$.
+Instead, issue resolution systems are evaluated by running acceptance tests $t_{\textrm{gold}}$ on $c_{\textrm{new}}$.
+In SWE-bench, $t_{\textrm{gold}}$ is mined from issue-resolving pull requests (Jimenez et al., 2024).
+Since most open issues lack acceptance tests, $t_{\textrm{gold}}$ is not yet available during issue resolution — neither in practice nor in the benchmark.
+Therefore, several issue resolution systems instead generate surrogate tests $t_{\textrm{gen}}$, based only on $d_{\textrm{issue}}$ and $c_{\textrm{old}}$.
+Since generated tests $t_{\textrm{gen}}$ are imperfect, over-relying on them risks test overfitting, where $c_{\textrm{new}}$ passes $t_{\textrm{gen}}$ but fails the hidden $t_{\textrm{gold}}$.
+issue description $d_{\textrm{issue}}$
+1JSONField values are displayed as dict when readonly in the admin.
+2For example, {"foo": "bar"} would be displayed as {’foo’: bar’}.
+code patch $c_{\textrm{new}}$ without refinement
+1+ elif isinstance(field, JSONField):
+2+ return field.prepare_value(value)
+observed generated test $t_{\textrm{gen}}$
+1value, disp = {"foo": "bar"}, {"foo": "bar"}’
+2self.assertEqual(
+3 display_for_field(value, JSONField(), empty), disp)
+code patch $c^{\prime}_{\textrm{new}}$ with refinement
+1+ elif isinstance(field, JSONField):
+2+ try:
+3+ return json.dumps(value)
+4+ except (TypeError, ValueError):
+5+ return empty_value_display
+hidden golden test $t_{\textrm{gold}}$
+1for value, disp in [
+2 ({a’: {b’: c’}}, {"a": {"b": "c"}}’),
+3 ([a’, b’], ["a", "b"]’),
+4 (a’, "a"’),
+5 ({(a’, b’): c’}, "{(’a’, ’b’): ’c’}"), # Invalid JSON.
+6]:
+7 self.assertEqual(
+8 display_for_field(value, JSONField(), empty), disp)
+Figure 1. Test overfitting example on django__django-12308.
+Figure 1 shows a concrete example.
+Based on the issue description $d_{\textrm{issue}}$, an issue resolution system generates initial new code $c_{\textrm{new}}$ and a reproduction test generator generates a test $t_{\textrm{gen}}$.
+However, $c_{\textrm{new}}$ fails on $t_{\textrm{gen}}$, triggering code refinement based on observed test $t_{\textrm{gen}}$.
+The refined code $c^{\prime}_{\textrm{new}}$ passes $t_{\textrm{gen}}$ but overfits to it, since it fails other cases covered by the hidden golden test $t_{\textrm{gold}}$.
+This paper empirically investigates test overfitting in issue resolution using popular large-scale open-source repositories and issues.
+It was inspired by two earlier studies.
+The paper “Is the cure worse than the disease?” (Smith et al., 2015) demonstrated that test overfitting is a serious problem.
+But unlike our paper, it predated large language models (LLMs) and focused on student projects.
+The “Inference scaling FLaws” (Stroebl et al., 2024) paper found test overfitting to be a limiting factor for inference scaling in LLM-based code generation.
+But unlike our paper, it did not study repository-level tasks, nor did it explore test-based code refinement.
+Overall, while these papers highlighted the problem in adjacent settings, to our knowledge, ours is the first to explore it in the setting of LLM-based issue resolution.
+Our empirical study answers three research questions:
+  • RQ1.
+    Do LLM-generated code patches $c_{\textrm{new}}$ overfit to tests $t_{\textrm{gen}}$ generated using the same LLM?
+  • RQ2.
+    How does refining code $c_{\textrm{new}}$ based on observed tests $t_{\textrm{gen}}$ affect test overfitting?
+  • RQ3.
+    If the hidden tests $t_{\textrm{gold}}$ were revealed, would refining code based on them break other functionality?
+Our experiments use two LLMs, Claude-3.7 Sonnet and GPT-4o.
+We use strong existing LLM-based systems as a starting point, obtaining the candidate code $c_{\textrm{new}}$ from Agentless (Xia et al., 2025) and the surrogate test $t_{\textrm{gen}}$ from e-Otter++ (Ahmed et al., 2025).
+We extend these systems with a test-based code refinement loop that is representative of those found in recent work (Ehrlich et al., 2025; Li et al., 2025c, b, a).
+The experiments are based on the TDD-bench Verified (Ahmed et al., 2024) benchmark for test-driven development, which is derived from SWE-bench (Jimenez et al., 2024).
+TDD-bench Verified comprises challenging real-world Python issue resolution tasks.
 This paper makes the following contributions.
-  • It measures the degree of test overfitting by LLMs in repository-level APR tasks.
-  • It describes techniques to reduce overfitting and reports how much they help overfitting and hurt performance.
-  • It also includes a limit study of how much an oracle with access to the golden black-box tests could help performance.
-Our paper title is inspired by Smith et al.
-(smith_et_al_2015).
-If the disease is code bugs and the cure is APR with white-box tests, Smith et al.
-found that the cure was indeed worse than the disease: APR decreased performance on black-box tests.
-Our findings are less drastic: on average across all benchmark instances, APR with white-box tests improves performance on black-box tests.
-But there is also still test overfitting to beware of—the cure can still cause harm.
+  • It measures the degree of test overfitting by LLMs in repository-level issue resolution tasks.
+  • It describes techniques to reduce overfitting and reports how they impact both overfitting and performance.
+  • It also includes a limit study of how much an oracle with access to the hidden golden tests could help performance.
+This paper demonstrates that test overfitting is a clear problem in modern LLM-based issue resolution systems.
+And unfortunately, while attempts to mitigate this overfitting do work to some extent, they also reduce effectiveness on the original task of issue resolution.
+Even if golden tests were available, refining code to pass them can still break other functionality.
+Overall, we caution against over-reliance on tests during code generation and hope to inspire future work in reducing test overfitting.
 2. Background and Related Work
-Test overfitting happens when an APR workflow generates code that passes white-box tests but fails held-out black-box tests (smith_et_al_2015).
+Test overfitting happens when an system generates code that passes observed tests but fails held-out hidden tests (Smith et al., 2015).
 Test overfitting is similar to overfitting a model to its training data, where the model performs worse on held-out test data; but note that in this work, we use only pre-trained LLMs without fine-tuning.
-Instead, test overfitting means that generated repaired code is too narrow to generalize to held-out black-box tests.
+Instead, test overfitting means that generated code $c_{\textrm{new}}$ is too narrow to generalize to held-out tests $t_{\textrm{gold}}$.
 Stroebl et al.
-report that this happens when a loop repeatedly generates code for the HumanEval or MBPP benchmarks until finding code that passes white-box tests (stroebl_kapoor_narayanan_2024).
-Unlike Stroebl et al., we experiment with repository-level benchmarks and we also explore workflows that improve code based on tests.
-SWE-bench is an APR benchmark where the input comprises an issue and the original code of an entire repository before issue resolution and the output is repaired code that must pass a held-out test set (jimenez_et_al_2024).
-Those held-out tests are unavailable, hidden from the APR workflow.
-SWE-bench Verified is a subset of 500 high-quality SWE-bench instances, targeted by most LLM-based APR workflows (chowdhury_et_al_2024).
-SWT-bench (mundler_et_al_2024) and TDD-Bench Verified (ahmed_et_al_2024) are reproduction test benchmarks derived from SWE-bench and SWE-Bench Verified.
-A generated output test must fail on the original code to reproduce the issue, and must pass on the hidden ground-truth code patch to confirm that it resolves the issue (mundler_et_al_2024).
-Several LLM-based workflows leverage white-box tests.
-CodeT generates many code candidates and many candidate white-box tests, then uses a dual execution agreement to pick a single output code (chen_et_al_2023).
-Agentless generates up to 40 candidate codes for SWE-bench and generates a white-box reproduction test, then uses the test to filter the candidate codes as part of code selection (xia_et_al_2025).
-Both S* (li_et_al_2025_sstar) and CodeMonkeys (ehrlich_et_al_2025) also generate both candidate codes and candidate tests, followed by a loop where the LLM improves the code based on the test and vice versa.
+report that this happens when a loop repeatedly generates code for the HumanEval or MBPP benchmarks until finding code that passes observed tests (Stroebl et al., 2024).
+Unlike Stroebl et al., we experiment with repository-level benchmarks and we also explore test-based code refinement loops.
+SWE-bench is a benchmark where the input comprises an issue description $d_{\textrm{issue}}$ and the original code $c_{\textrm{old}}$ of an entire repository before issue resolution and the output is repaired code $c_{\textrm{new}}$ (Jimenez et al., 2024).
+Solutions are evaluated on whether $c_{\textrm{new}}$ passes a held-out test set $t_{\textrm{gold}}$.
+To emulate the real-world situation where issues do not come with readily executable tests, the benchmark does not make those held-out tests available to the issue resolution system.
+SWE-bench Verified is a subset of 500 high-quality SWE-bench instances, targeted by most issue resolution systems (Chowdhury et al., 2024).
+SWT-bench (Mündler et al., 2024) and TDD-Bench Verified (Ahmed et al., 2024) are reproduction test benchmarks derived from SWE-bench and SWE-Bench Verified.
+Several issue resolution systems leverage generated tests $t_{\textrm{gen}}$.
+CodeT generates many code candidates and many candidate tests, then uses a dual execution agreement to pick a single output code (Chen et al., 2023).
+Agentless generates up to 40 candidate codes for SWE-bench and generates a reproduction test, then uses the test to filter the candidate codes as part of code selection (Xia et al., 2025).
+S* (Li et al., 2025a), CodeMonkeys (Ehrlich et al., 2025), and InfCode (Li et al., 2025c) also generate both candidate codes and candidate tests, followed by a loop where the LLM improves the code based on the test and vice versa.
 These papers might suffer from test overfitting but do not study it.
 Test overfitting is related to reward hacking, where a reward is a reinforcement-learning (RL) objective, and a hack finds a loop-hole for improving the reward while circumventing its intention.
-CURE uses RL to co-evolve an LLM to generate both code and white-box tests, and might be susceptible to test overfitting and reward hacking (wang_et_al_2025).
+CURE uses RL to co-evolve an LLM to generate both code and white-box tests, and might be susceptible to test overfitting and reward hacking (Wang et al., 2025).
 Baker et al.
-found that RL to improve frontier LLMs for repository-level APR tasks can cause the reward hack of disabling tests instead of repairing code (baker_et_al_2025).
+found that RL to improve frontier LLMs for repository-level coding tasks can cause the reward hack of disabling tests instead of improving code (Baker et al., 2025).
 3. Methodology
-Let $c_{\mathrm{old}}$ be the original code in the repository before a given issue is fixed.
-This work uses Agentless (xia_et_al_2025) to generate new code $c_{\mathrm{new}}$ for APR and e-Otter++ (ahmed_et_al_2025_eotter) to generate a white-box test $t$ for issue reproduction.
-Being generated by LLMs, both $c_{\mathrm{new}}$ and $t$ are possibly imperfect.
-We run the generated test $t$ on the generated code $c_{\mathrm{new}}$.
-If it fails, we use a test-based code improvement approach to generate a better $c_{\mathrm{new}}$ that passes the (possibly also improved) white-box test $t$.
-This section briefly discusses the generation of the initial $c_{\mathrm{new}}$ and $t$ and our code improvement approach.
-3.1. Code Patch Generation: Agentless
-Agentless performs hierarchical localization followed by repair and patch validation (xia_et_al_2025).
-Localization starts by identifying suspicious files using both LLM prompting and embedding-based retrieval, filtering out irrelevant files.
-It then narrows down to relevant classes or functions using a concise skeleton representation of files.
-Finally, it pinpoints exact edit locations at the line level.
-Repair generates multiple candidates for new code $c_{\textrm{new}}$ in a lightweight search-replace diff format rather than rewriting entire code blocks, which minimizes hallucinations and enhances cost-efficiency.
-Patch validation automatically generates reproduction tests $t$ (when none are provided) and combines them with existing regression tests in the repository to select a single $c_{\textrm{new}}$ via majority voting.
-This modular structure allows Agentless to remain interpretable, easily debuggable, and effective without relying on complex environment interaction or planning mechanisms typical of agents.
-With the Claude-3.7-Sonnet model, Agentless achieves a 50.8% issue resolution rate on SWE-bench Verified.
-3.2. Reproduction Test Generation: e-Otter++
-e-Otter++ (ahmed_et_al_2025_eotter)is a system designed to generate reproduction tests $t$ from software issue descriptions, such as bug reports or feature requests, before the corresponding code patch exists.
-It supports test-driven development workflows and helps validate code patches produced by APR workflows.
-The system augments LLM outputs with checks and repairs, such as linter checks and fixing import statements.
-It also employs a self-reflective action planner that iteratively decides which parts of the code to ‘read’ and where and how to write or modify test code.
-After generating candidate tests, e-Otter++ obtains execution feedback, such as runtime errors or successes, and uses this information to refine prompts and guide test selection.
-Additionally, e-Otter++ leverages issue morphing–transforming a bug report into multiple variants to generate diverse tests.
-It then uses a ranker that combines running on imperfect surrogate code patches and coverage to select a single top-ranked test $t$.
-As a result, e-Otter++ achieves a fail-to-pass success rate of 63% on the TDD-Bench Verified benchmark using Claude-3.7-Sonnet.
-3.3. Test-based Code Refinement
-Figure 1. Overview of test-based code improvement.
+We assume a simple three-step pipeline for issue resolution.
+  1. (1)
+    Run an off-the-shelf issue resolver (Agentless (Xia et al., 2025)) on $d_{\textrm{issue}}$ and original code $c_{\textrm{old}}$ to obtain initial candidate code $c_{\textrm{new}}$.
+  2. (2)
+    Run an off-the-shelf reproduction test generator (e-Otter++ (Ahmed et al., 2025)) on $d_{\textrm{issue}}$ and $c_{\textrm{old}}$ to obtain an initial candidate test $t_{\textrm{gen}}$.
+  3. (3)
+    Run test-based code refinement (described below) on the initial $c_{\textrm{new}}$ and $t_{\textrm{gen}}$ to obtain final candidate code $c^{\prime}_{\textrm{new}}$.
+After that, we can quantify test overfitting by checking which code ($c_{\textrm{new}}$ or $c^{\prime}_{\textrm{new}}$) passes which tests ($t_{\textrm{gen}}$, $t_{\textrm{gold}}$, or $t_{\textrm{old}}$).
+Since Agentless (Xia et al., 2025) and e-Otter++ (Ahmed et al., 2025) have been described in prior work, here we only briefly discuss characteristics salient to our paper.
+First, both Agentless and e-Otter++ internally employ some parallel inference scaling.
+In our pipeline, we only take the single final result selected by each system and feed it into our own test-based code refinement stage, which can be viewed as a form of sequential inference scaling.
+Second, both Agentless and e-Otter++ internally make use of regression tests $t_{\textrm{old}}\subset c_{\textrm{old}}$ from the original code of the repository.
+This is a realistic reflection of how a human software engineer might consult regression tests when tackling an issue.
+At the same time, it only yields limited benefit, since existing tests $t_{\textrm{old}}$ tend not to test the open issue $d_{\textrm{issue}}$.
+Figure 2. Overview of test-based code refinement.
 Figure 2 shows our test-based code refinement approach.
-We start with the initial candidate code $c_{\textrm{new}}$ from Agentless and the initial candidate test $t$ from e-Otter++.
+It starts with the initial candidate code $c_{\textrm{new}}$ from Agentless and the initial candidate test $t_{\textrm{gen}}$ from e-Otter++.
 We execute the test on $c_{\textrm{old}}$ and $c_{\textrm{new}}$.
 If the test goes from fail-to-pass, we immediately stop and declare success.
-Otherwise, we collect the execution logs from both test executions, the focal functions modified by $c_{\textrm{new}}$, and the test function $t$ and provide them to an LLM-based critic.
+Otherwise, we collect the execution logs from both test executions, the focal functions modified by $c_{\textrm{new}}$, and the test function $t_{\textrm{gen}}$ and provide them to an LLM-based critic.
 We ask the critic which part to modify (focal or test).
 Inspired by e-Otter++, we also collect additional information such as the buggy line, relevant issue line, and lookup function (the function the model wants to see to fix it).
-Then, we ask the LLM to modify the focal or test, exposing the counterpart (e.g., for focal, the white-box test) and additional information.
-After updating the focal or test, if the test is still not fail-to-pass, we need to decide whether to stick with our initial focal or test or update it with the recently generated focal or test, even though the test is not fail-to-pass.
-To do this, we use a reward function to guide our decision: $\mathit{Reward}(c_{\mathit{old}},c_{\mathit{new}},t)\,=$
-$\frac{1}{3}\mathit{isFail}(t,c_{\mathrm{old}})+\frac{1}{3}\mathit{isPass}(t,c_{\mathrm{new}})+\frac{1}{3}\mathit{coverage}(c_{\mathit{old}},c_{\mathit{new}},t)$
+Then, we ask the LLM to modify the focal code or test, exposing the counterpart (e.g., for focal, the test) and additional information.
+After updating the code $c_{\textrm{new}}$ or test $t_{\textrm{gen}}$, if the test is still not fail-to-pass, we need to decide whether to stick with our initial code or test or update it with the recently generated code or test, even though the test is not fail-to-pass.
+To do this, we use a reward function to guide our decision: $\mathit{Reward}(c_{\mathit{old}},c_{\mathit{new}},t_{\textrm{gen}})\,=$
+$\frac{1}{3}\mathit{isFail}(t,c_{\mathrm{old}})+\frac{1}{3}\mathit{isPass}(t,c_{\mathrm{new}})+\frac{1}{3}\mathit{coverage}(c_{\mathit{old}},c_{\mathit{new}},t_{\textrm{gen}})$
 In the reward function, the first two components are binary, but coverage can vary from 0 to 1.
-Since we know both the patch and the test, we can compute how many deleted lines in $c_{\mathrm{old}}$ and how many added lines in $c_{\mathrm{new}}$ have been covered by the test.
+Since we know both the code patch and the test, we can compute how many deleted lines in $c_{\mathrm{old}}$ and how many added lines in $c_{\mathrm{new}}$ have been covered by the test.
 We divide the covered lines by the total number of added or deleted lines to compute coverage.
-If our new code+test pair achieves a higher score than the previous one, we replace it with the new one.
-We repeat the improvement loop for at most 15 iterations if the white-box test does not change from fail to pass.
+If our new $\langle c_{\mathit{new}},t_{\textrm{gen}}\rangle$ pair achieves a higher reward score than the previous one, we replace it with the new one.
+We repeat the improvement loop for at most 15 iterations if the observed test $t_{\textrm{gen}}$ does not change from fail to pass.
 4. Results
-4.1. Expermental Setup
 The experiments in this paper are based on Agentless [^1] https://github.com/OpenAutoCoder/Agentless-generated code patches and e-Otter++ [^2] https://zenodo.org/records/16755224-generated tests.
@@
 We evaluated on all 449 instances.
-The experiments use two frontier models: Claude-3.7-Sonnet and GPT-4o.
 Our primary metric is test overfitting rate.
-A code patch is overfitting tests if it passes white-box tests but fails black-box tests.
-Dividing the number of instances where that happes by the total number of instances yields the overfitting rate.
+A code patch $c_{\textrm{new}}$ is overfitting tests if it passes $t_{\textrm{new}}$ but fails $t_{\textrm{gold}}$ or $t_{\textrm{old}}$.
+Dividing the number of instances where that happens by the total instances yields the overfitting rate.
 Figure 3. How severe is test overfitting? (Claude-3.7-Sonnet)
-Table 1. Test overfitting rate of automatic program repair.
+Table 1. Test overfitting rate of automated issue resolution.
 Model Setup
@@
 w/ Refinement 198 127 71 35.9
-4.2. LLMs Overfit on White-box Tests (RQ1)
-To answer this research question, we run the e-Otter++ generated white-box tests on Agentless patches and see how many of them go from fail-to-pass.
-Figure 2 shows that with the Claude-3.7-Sonnet model, 229 tests actually pass on the initial patches.
-Does this mean that the black-box golden tests will pass on these patches?
+4.1. Overfitting without Code Refinement (RQ1)
+To quantify overfitting without refinement, we run the e-Otter++ generated tests $t_{\textrm{gen}}$ on Agentless code patches $c_{\textrm{new}}$ and see how many pass.
+Figure 3 shows that with the Claude-3.7-Sonnet model, for 229 instances, the initial code passes the generated tests.
+Does this mean that the hidden golden tests $t_{\textrm{gold}}$ will pass on these patches?
 The answer is no.
-There are 50 instances where the black-box test could not pass, signaling potential overfitting (50/229=21.8%).
+There are 50 instances where $t_{\textrm{gold}}$ failed, rendering them unresolved according to the benchmark.
+That means the overfitting rate among the 229 instances is $\frac{50}{229}=21.8\%$.
 For the GPT-4o model, the overfitting rate is 33.0%, even higher than the Claude-3.7-Sonnet model (see Table 1).
@@
 What about using tests to improve patches at generation time?
-Using our test-based code refinement technique, we attempt to generate an improved code+test pair for each of the 220 code+test pairs where the tests did not initially go from fail to pass.
-Refinement succeeded in generating 22 pairs where the white-box test goes from fail to pass.
-Unfortunately, 14 out of the 22 fail black-box testing, increasing the overfitting.
-If we include these in our existing overfitted samples, the overfitting rate increases from 21.8% to 25.5% (Table 1).
+We apply the test-based code refinement loop from Section 3 to each of the 220 instances from Figure 3 where the initial code $c_{\textrm{new}}$ failed the observed generated tests $t_{\textrm{gen}}$.
+This yielded 22 instances where the refined code $c^{\prime}_{\textrm{new}}$ passed the generated tests.
+Unfortunately, 14 out of the 22 fail the hidden tests $t_{\textrm{gold}}$, increasing the overfitting.
+If we include these in our existing overfitted samples, the overfitting rate increases to $\frac{50\,+\,14}{229+22}=\frac{64}{251}=25.5\%$ (see Table 1).
 GPT-4o experiences a similar increase in overfitting, indicating that it is difficult to improve code patches by exposing tests to LLMs.
-In the above experiments, the refinement loop exposed the white-box tests and additional information to the LLMs.
-One idea to reduce overfitting might be to hide the test and additional information, so the LLMs may not be able to trick the test.
-Indeed, hiding the test causes a drop in agreement between code patches and white-box tests.
-However, the number of instances passing black-box tests also goes down a bit.
+In the above experiments, the refinement loop exposed the generated tests $t_{\textrm{gen}}$ and additional information to the LLMs.
+One idea to reduce overfitting might be to hide the tests and additional information, so the LLMs may not be able to trick the tests.
+Indeed, hiding the generated tests and revealing merely a pass/fail flag to the LLM causes a drop in agreement between code patches and generated tests.
+However, the number of instances passing the hidden tests $t_{\textrm{gold}}$ also goes down a bit.
 If we remove the additional information, it becomes more difficult for LLMs to overfit the patch.
@@
 Note that this will not have any impact on instances that were already overfitted before the refinement.
-Table 2. Test overfitting rate of repaired samples.
+Table 2. Test overfitting rate of repaired samples, and effect over hiding information to mitigate overfitting.
 Model
@@
 Overfitting
-Claude-3.7 Sonnet 22 7 15 68.2
+Claude-3.7 Sonnet 22 8 14 63.6
 18 5 13 72.2
@@
 16 5 11 68.8
-Table 2 shows that exposing tests and prompt augmentation resolved 7 samples.
-Do all these 7 samples help us resolve additional instances?
+Table 2 shows that exposing tests and prompt augmentation resolved 8 samples.
+Do all these 8 samples help us resolve additional instances?
 The answer is no.
-The initial code patch for one instance was already passing on the black-box test.
-And besides these 7, there were 2 instances that went from resolved to non-resolved.
-So the net gain is +4 instead of +7.
+The initial code patch for one instance was already passing on the golden test.
+And besides these 8, there were 2 instances that went from resolved to non-resolved.
+So the net gain is +5 instead of +8.
 That indicates the actual gain from test-based refinement is less than the apparent gain.
-4.4. Impact of Revealing Black-box Tests (RQ3)
-Table 3 shows how much overfitting occurs when using some of the golden tests for code selection and repair.
-For this experiment, we divide black-box tests into two groups: fail-to-pass reproduction tests and pass-to-pass regression tests.
-We only reveal the reproduction tests (so they are no longer black-box) but continue to hide the regression tests.
-This is a limit study, because in practice, all black-box tests are hidden.
-The overfitting rate, which means passing now-revealed reproduction tests but failing still-hidden regression tests, is low (13/223=5.8% and 22/194=11.3% with Claude and GPT-4o, respectively).
-Apart from overfitting, this also shows how well the model performs when some black-box tests are exposed.
+4.3. Impact of Revealing Golden Tests (RQ3)
+RQ3 asks what would happen if the test-based code refinement would not use imperfect generated tests $t_{\textrm{gen}}$ but rather golden tests $t_{\textrm{gold}}$.
+This is a limit study, because in practice, issues do not usually come with oracular executable acceptance tests.
+Table 3 shows how much overfitting occurs in this setting.
+Here, the final determination of whether an instance is resolved is based on passing both golden tests $t_{\textrm{gold}}$ and a subset of regression tests $t_{\textrm{old}}$.
+(While issue-resolution systems such as Agentless have access to the full regression testing suite contained in the original code $c_{\textrm{old}}$, they do not know which subset of those tests are expected to pass after issue resolution.) The overfitting rate, which means passing now-revealed golden tests $t_{\textrm{gold}}$ but failing the relevant subset of regression tests $t_{\textrm{old}}$, is low (13/223=5.8% and 22/194=11.3% with Claude and GPT-4o, respectively).
+Apart from overfitting, this also shows how well the model performs when golden tests $t_{\textrm{gold}}$ are exposed.
 When the model passes the golden reproduction tests, the probability that the issue is resolved is high, but not 100%, because regression tests can still fail.
-The improvement from using black-box tests in issue resolution is modest: the number of resolved instances with the Claude model goes up from 229 to 243, which is around a 3% increase.
-Overfitting in this setup is much less since we are evaluating against the golden tests–however, the golden tests themselves are really just a proxy for “correctness”, against which we might still be overfitting (but we do not have a good way to measure this).
+The improvement from using golden tests in issue resolution is modest: the number of resolved instances with the Claude model goes up from 229 to 243, around a 3% increase.
+This experiment emulates a situation where developers follow test-driven development and write tests prior to generating $c_{\textrm{new}}$.
+Unfortunately, the gains are limited, highlighting that current issue resolution systems often miss the bigger picture.
 Table 3. Impact of revealing golden test on issue resolution.
@@
 5. Discussion
-5.1. LLMs’ Preference on Focal and Test
+LLMs’ Preference on Focal and Test
 Test-based code refinement gives LLMs a choice to modify both the code (focal function) and/or the test.
@@
 Figure 4. Coverage of the unbiased and overfitted patches.
-5.2. Coverage for Overfitted Sample
-Figure 3 compares the coverage between unbiased code patches (both white-box and black-box tests pass) and overfitted patches for Claude-3.7.
-Note that the tests are generated by e-Otter++ and the code patches are from Agentless.
+Coverage for Overfitted Sample
+Figure 4 compares the coverage between unbiased code patches (both $t_{\textrm{gen}}$ and $t_{\textrm{gold}}$ pass) and overfitted patches for Claude-3.7.
 The median for unbiased patches is 1, while for overfitted patches it is less than 0.8.
@@
 0.9).
-These results indicate that even though both unbiased and overfitted patches pass the white-box test, overfitted patches have lower coverage.
-5.3. What Could Happen with Overfitting?
-Frontier LLMs are capable of tricking the test and changing the code to narrowly passes that particular test.
-For example, instance id “astropy__astropy-14598” complains about an inconsistency in double and single quotes.
-The issue writer notes that while dealing with a null string, the double quotes become single quotes.
-To trick the white-box test, LLMs modify their initial solution (see Figure 4(a)) to a solution where “replace” is invoked three times just to pass the white-box test (see Figure 4(b)), which is clearly an indication of LLMs overfiting.
-We have even seen LLMs using reflection to access private methods to pass the white-box test, creating a security risk.
-(a) Initial code patch
-(b) Overfitted refined code patch
-Figure 4. Example where test-based refinement overfits.
+These results indicate that even though both unbiased and overfitted patches pass observed tests $t_{\textrm{gen}}$, overfitted patches have lower coverage.
 6. Threats to Validity
-One of the main limitations of our work is that the benchmarks include only Python repositories.
-Therefore, our findings may not generalize to other programming languages.
+Since the benchmarks include only Python repositories, our findings may not generalize to other programming languages.
 We experimented with two models but did not investigate model contamination or memorization.
@@
 We share the initial and modified code patches at this link: https://zenodo.org/records/17227319
-8. Future Plans
-This paper demonstrates that overfitting exists and shows attempts to minimize it in a test-based code repair system.
-However, we have not proposed any methods to address overfitting in cases where the initial white-box test changes from fail to pass on overfitted patches.
-This is an important problem because the majority of overfitted patches come from this category.
-In the future, we will try to design an approach to automatically detect these overfitted patches and propose a solution to address the issue.

### v3

@@
 LLMs, SWE Patches, Reproduction Tests
+journalyear: 2026copyright: ccconference: 34th ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering; July 05–09, 2026; Montreal, QC, Canadabooktitle: 34th ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering (FSE Companion ’26), July 05–09, 2026, Montreal, QC, Canadadoi: 10.1145/3803437.3805574isbn: 979-8-4007-2636-1/2026/07ccs: Software and its engineering Automatic programming
 1. Introduction
@@
 Our experiments use two LLMs, Claude-3.7 Sonnet and GPT-4o.
-We use strong existing LLM-based systems as a starting point, obtaining the candidate code $c_{\textrm{new}}$ from Agentless (Xia et al., 2025) and the surrogate test $t_{\textrm{gen}}$ from e-Otter++ (Ahmed et al., 2025).
+We use strong existing LLM-based systems as a starting point, obtaining the candidate code $c_{\textrm{new}}$ from Agentless (Xia et al., 2025) and the surrogate test $t_{\textrm{gen}}$ from e-Otter++ (Ahmed et al., 2026).
 We extend these systems with a test-based code refinement loop that is representative of those found in recent work (Ehrlich et al., 2025; Li et al., 2025c, b, a).
@@
   2. (2)
-    Run an off-the-shelf reproduction test generator (e-Otter++ (Ahmed et al., 2025)) on $d_{\textrm{issue}}$ and $c_{\textrm{old}}$ to obtain an initial candidate test $t_{\textrm{gen}}$.
+    Run an off-the-shelf reproduction test generator (e-Otter++ (Ahmed et al., 2026)) on $d_{\textrm{issue}}$ and $c_{\textrm{old}}$ to obtain an initial candidate test $t_{\textrm{gen}}$.
   3. (3)
@@
 After that, we can quantify test overfitting by checking which code ($c_{\textrm{new}}$ or $c^{\prime}_{\textrm{new}}$) passes which tests ($t_{\textrm{gen}}$, $t_{\textrm{gold}}$, or $t_{\textrm{old}}$).
-Since Agentless (Xia et al., 2025) and e-Otter++ (Ahmed et al., 2025) have been described in prior work, here we only briefly discuss characteristics salient to our paper.
+Since Agentless (Xia et al., 2025) and e-Otter++ (Ahmed et al., 2026) have been described in prior work, here we only briefly discuss characteristics salient to our paper.
 First, both Agentless and e-Otter++ internally employ some parallel inference scaling.
@@
 Our primary metric is test overfitting rate.
-A code patch $c_{\textrm{new}}$ is overfitting tests if it passes $t_{\textrm{new}}$ but fails $t_{\textrm{gold}}$ or $t_{\textrm{old}}$.
+A code patch $c_{\textrm{new}}$ is overfitting tests if it passes $t_{\textrm{gen}}$ but fails $t_{\textrm{gold}}$ or $t_{\textrm{old}}$.
 Dividing the number of instances where that happens by the total instances yields the overfitting rate.
@@
 For the GPT-4o model, the overfitting rate is 33.0%, even higher than the Claude-3.7-Sonnet model (see Table 1).
+Note that a test can overfit due to weaknesses in both the code patches and the tests.
+However, we did not explicitly differentiate between the two, as doing so would require extensive qualitative studies, which are difficult to scale.
 4.2. Overfitting with Code Refinement (RQ2)
@@
 That indicates the actual gain from test-based refinement is less than the apparent gain.
+The refinement reward’s coverage term may structurally push patches toward fitting $t_{\textrm{gen}}$’s execution path.
+However, it is also believed that tests with higher coverage are qualitatively better than other tests.
 4.3. Impact of Revealing Golden Tests (RQ3)
@@
 Here, the final determination of whether an instance is resolved is based on passing both golden tests $t_{\textrm{gold}}$ and a subset of regression tests $t_{\textrm{old}}$.
-(While issue-resolution systems such as Agentless have access to the full regression testing suite contained in the original code $c_{\textrm{old}}$, they do not know which subset of those tests are expected to pass after issue resolution.) The overfitting rate, which means passing now-revealed golden tests $t_{\textrm{gold}}$ but failing the relevant subset of regression tests $t_{\textrm{old}}$, is low (13/223=5.8% and 22/194=11.3% with Claude and GPT-4o, respectively).
+(While issue-resolution systems such as Agentless have access to the full regression testing suite contained in the original code $c_{\textrm{old}}$, they do not know which subset of those tests are expected to pass after issue resolution (Chen et al., 2026).) The overfitting rate, which means passing now-revealed golden tests $t_{\textrm{gold}}$ but failing the relevant subset of regression tests $t_{\textrm{old}}$, is low (13/223=5.8% and 22/194=11.3% with Claude and GPT-4o, respectively).
 Apart from overfitting, this also shows how well the model performs when golden tests $t_{\textrm{gold}}$ are exposed.
@@
 Table 4 shows that for instances where refinement succeeds, the model mostly elects to modify the focal function instead of the test.
-For each scenario, the number of test updates is very low compared to focal updates.
-Also, there are few cases where we see the LLMs try to fix both the focal function and the test.
+For each scenario, the number of test updates is low compared to focal updates.
+Also, there are few cases where the LLMs tries to fix both the focal function and the test.
 That indicates the model has a preference for focal modification and believes that the tests are mostly perfect.
@@
 (Gold) (Gold) 28 20 2 6
-Figure 4. Coverage of the unbiased and overfitted patches.
 Coverage for Overfitted Sample
-Figure 4 compares the coverage between unbiased code patches (both $t_{\textrm{gen}}$ and $t_{\textrm{gold}}$ pass) and overfitted patches for Claude-3.7.
+We compare the coverage between unbiased code patches (both $t_{\textrm{gen}}$ and $t_{\textrm{gold}}$ pass) and overfitted patches for Claude-3.7.
 The median for unbiased patches is 1, while for overfitted patches it is less than 0.8.
@@
 These results indicate that even though both unbiased and overfitted patches pass observed tests $t_{\textrm{gen}}$, overfitted patches have lower coverage.
+Execution coverage or static analysis might be used as a warning against overfitting early during patch generation, without relying on hidden golden tests.
+We leave this for future work.
 6. Threats to Validity
@@
 Therefore, using tests for patch generation needs more deliberation.
-We share the initial and modified code patches at this link: https://zenodo.org/records/17227319
+We share the initial and modified code patches at this link: https://doi.org/10.5281/zenodo.17227318